Circular No. 6469

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996as IN ANONYMOUS GALAXY B. P. Schmidt, Mount Stromlo and Siding Spring Observatories, re-

ports the discovery of a supernova (mag ∼ 17.5) by R. O. Evans on a U.K. Schmidt film taken by M. Drinkwater on Sept. 3 as part of the U.K. Schmidt Supernova Search. SN 1996as is located at α = 21^{h}05^{m}36^{s}.3, δ = −52^{o}22′41′′.2 (equinox 2000.0), which is ∼ 15′′ northwest of the galaxy’s nucleus. SN 1996as was not visible on a similar film taken on Aug. 14.

S. Benetti and M. Turatto, European Southern Observatory (ESO), obtained a spectrum (range 515–927 nm, resolution 1.1 nm) of SN 1996as on Sept. 6.02 UT at La Silla with the 2.2-m ESO/MPI telescope (+ EFOSC-2). The object is a type-II supernova with a blue (T_{BB} ∼ 10 400 K) continuum and broad emission lines of Hα (FWHM = 13 800 km/s) and He I (FWHM = 10 000 km/s). The spectrum’s continuum slope and line profiles resemble that of SN 1992H on day 13 after explosion (Clocchiatti et al. 1996, A.J. 111, 1286). The redshift of the galaxy, measured from narrow emissions of Hα and [N II] superimposed on the galaxy spectrum, is 0.036.

γ-RAY BURST 960720 T. Boller and W. Voges, Max-Planck-Institut für Extraterrestrische

Physik, Garching, report: “Within the 10′ error box centered on the SAX position of the July 20 γ-ray burst (IAUC 6467), we have found a faint x-ray source in the ROSAT all-sky survey data. The ROSAT source is centered at α = 17^{h}29^{m}34^{s}, δ = +49^{o}05′.8 (equinox 2000.0; estimated error 30′′ in 1σ radius), which is within 3′.7 of the SAX position. The PSPC countrate in the ROSAT hard-energy band (0.5–2.0 keV) was 0.007 ± 0.002 count/s during 1990 Aug. 22–28. This is the only ROSAT x-ray source within the SAX error circle. It is not clear if this source relates to the γ-ray burst.”

1996 PW I. P. Willams, Queen Mary and Westfield College, London; and M.

Cartwright and A. Fitzsimmons, Queen’s University, Belfast, report the following B, R, and B − R photometry of 1996 PW made at the 1.0-m Jacobus Kapteyn Telescope on La Palma: Sept. 2.0111 UT, –, 16.81, –; 2.0306, 18.12, –, +1.31; 3.0014, –, 17.15, –; 3.0174, 18.39, –, +1.24; 3.0361, –, 17.12, +1.27; 4.0174, –, 17.09, –; 4.0299, 18.33, –, +1.24.

1996 September 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green